# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

### v2

@@ 9.1 Details of Prompt for MPO
Prompt Filtering. We employ the prompt filtering approach proposed by Zhang et al. [51] to curate our dataset. This strategy comprises two pivotal steps:
  • Semantic-Based Filtering: Utilizing an existing scene graph parser [46], we evaluate the semantic richness of prompts by analyzing the number of subjective and objective relationships. Prompts with fewer than one meaningful relationship are filtered out to reduce noise data.
@@
Dataset Curation. We apply this filtering strategy to the combined training sets from the Pick-a-Pick and HPS datasets, which initially comprise 166,475 prompts. Following comprehensive filtering, we retain 63,165 high-quality prompts for subsequent Multiobjective Preference Optimization (MPO) training.
@@ 9.2 VisionReward for Text-to-Image Optimization
Evaluation Settings. We conducted both automatic and human evaluation on DrawBench [35]. Automatic evaluation includes multiple metrics such as human preference RMs, CLIP [32] and LAION-Aesthetic [36]. Human evaluation requires annotators to comprehensively evaluate two images and select the better one, with an option to indicate a tie if neither is preferred. Annotators are required to take different perspectives (such as composition, image quality, realism, safety, and image-text alignment) into account.
Experimental Results. Main results are demonstrated in  Tab. 16 and  Tab. 17. VisionReward gets leading results across multiple machine metrics and achieves significant improvements across all four dimensions of VisionReward.
@@ 9.3 More Results of MPO
Training Curve. Fig. 10 shows the variation of the dimensional scores during the MPO process with respect to the number of training samples. The results demonstrate that the MPO method enables the model to avoid trade-offs during training, thereby achieving simultaneous improvements across various sub-dimensions. In contrast, the DPO [43] method fails to achieve this level of concurrent enhancement.
@@
10 Details of Dimensions for Annotation
Tab. 18 and  Tab. 19 show the annotation taxonomy of images, while  Tab. 20 and  Tab. 21 show for videos.
Weight and Accuracy of Checklist. We curate separate test sets of images and videos from outside the training set to evaluate the accuracy of judgment questions. The test set comprises 1,209 images and 1,000 videos, respectively. We report the accuracy of judgment questions (Cf.  Tab. 22 and  Tab. 23 for text-to-image,  Tab. 24 and  Tab. 25 for text-to-video). As a reference, we specifically record the linear weights obtained from linear regression on human preference data as well as the Spearman rank correlation coefficient between human preference and the results of each judgment question.
@@ 10 Details of Dimensions for Annotation
Fidelity Detail realism realistic Are the image details realistic?
@@ 11 More Details of MonetBench
Image-MonetBench Construction. We first establish our dataset foundation by collecting 4,038 seed prompts through strategic sampling from established datasets (1,000 from ImageRewardDB [49], 1,000 from HPDv2 [49], and 2,038 from Pick-a-Pic [22]). Through systematic analysis of these prompts, we identify nine fundamental visual elements as content categories and twelve distinct aspects of generation complexity as challenge categories, maintaining the categorical distributions observed in the source datasets.
Video-MonetBench Construction. For video prompt evaluation, we initially sample 20,000 prompts from the VproM [44] dataset, which are filtered to 13,342 valid entries after removing duplicates and invalid content. Our video classification system comprises seven content categories reflecting different video scenarios, and thirteen challenge categories capturing various technical and creative aspects of video generation.
Prompt Generation and Filtering. To ensure benchmark quality and diversity, we employ ChatGLM [11] to generate 1,000 new prompts for each benchmark following the established category distributions. Each generated prompt undergoes a three-stage filtering process: (1) Rouge-L [26] similarity checking for textual diversity, (2) semantic filtering with a cosine similarity threshold of 0.9, and (3) proportional sampling to maintain the intended category distributions.
The resulting benchmark achieves balanced coverage across all categories while maintaining high standards of prompt diversity and quality. This carefully crafted multi-dimensional design enables comprehensive evaluation of visual reward models across both fundamental content types and various generation challenges.
@@
Details of Classification Proportions. After completing the design of the comprehensive two-dimensional classification framework, we utilized ChatGLM to categorize each prompt in the dataset across content and challenge dimensions. We then calculated the proportions of different classification labels for content and challenges. The content and challenge categories and their respective examples are summarized in Tables 26 and 27. Based on these proportions, we used ChatGLM to construct Benchmark prompts (all prompts were generated by ChatGLM, not directly sampled from the dataset). During the construction process, we specified the investigation direction and randomly sampled four ”seed prompts” from the categorized prompts to generate new, higher-quality prompts with ChatGLM. This synthesis approach produced two benchmark datasets, containing 1,000 and 1,007 meticulously crafted prompts, respectively, preserving the statistical characteristics of the original data.

### v3

@@ 9.1 Details of Prompt for MPO
We employ the prompt filtering approach proposed by UniFL (zhang2024uniflimprovestablediffusion) to curate our dataset. This strategy comprises two pivotal steps:
  • Semantic-Based Filtering: Utilizing an existing scene graph parser (8954449), we evaluate the semantic richness of prompts by analyzing the number of subjective and objective relationships. Prompts with fewer than one meaningful relationship are filtered out to reduce noise data.
@@ 9.2 VisionReward for Text-to-Image Optimization
Evaluation Settings. We conducted both automatic and human evaluation on DrawBench (saharia2022photorealistic). Automatic evaluation includes multiple metrics such as human preference RMs, CLIP (radford2021learning) and LAION-Aesthetic (schuhmann2022laion).
Experimental Results. Main results are demonstrated in  Table 13 and  Table 14. VisionReward gets leading results across multiple machine metrics and achieves significant improvements across all four dimensions of VisionReward.
@@ 9.3 More Results of MPO
Training Curve. Fig. 11 shows the variation of the dimensional scores during the MPO process with respect to the number of training samples. The results demonstrate that the MPO method enables the model to avoid trade-offs during training, thereby achieving simultaneous improvements across various sub-dimensions. In contrast, the DPO (wallace2024diffusion) method fails to achieve this level of concurrent enhancement.
Ablation Study: Different Strategy for MPO. In using the MPO strategy, we define the way in which $R^{i}$ dominates $R^{j}$ (given two images $x^{i}$ and $x^{j}$) to select pairs. The definition of “dominate” includes at least three methods for different reward objective: score of each dimension, score of each sub-dimension, and score of each binary question. To investigate the impact of different definitions of “dominate”, we conduct experiments based on CogVideoX-5B. Specifically, we employ three different strategies, setting the threshold of the total score to 0.6, 0.5, and 0.4 respectively, ensuring that the number of pairs obtained through all three strategies is 5k. We use a batch size of 64, a learning rate of 2e-6, the DPO parameter $\beta$ of 500, and the training steps of 300. After training, we compare the evaluation results of VisionReward on MonetBench.  Table 15 shows that using the score for each dimension as objective yields the best results.
Methods Baseline Dimension Sub-dimension Question
Reward 4.303 4.573 4.539 4.514
@@
10 Details of Fine-Grained Questions
Table 16 and  Table 17 show the annotation taxonomy of images, while  Table 18 and  Table 19 show for videos.
11 More Results of Fine-Grained Design
Weight and Accuracy of Checklist. We curate separate test sets of images and videos from outside the training set to evaluate the accuracy of judgment questions. The test set comprises 1,209 images and 1,000 videos, respectively. We report the accuracy of judgment questions (Cf.  Table 20 and  Table 21 for text-to-image,  Table 22 and  Table 23 for text-to-video). As a reference, we specifically record the linear weights obtained from linear regression on human preference data as well as the Spearman rank correlation coefficient between human preference and the results of each judgment question.
@@ 11 More Results of Fine-Grained Design
Fidelity Detail reality realistic Are the image details realistic?
@@
12 Limitations
This section outlines several limitations of the VisionReward framework that emerged during its development.
Supported Video Frame Rate and Length. While training datasets of VisionReward contain videos up to 6 seconds in duration at 4 frames per second (fps), this may be insufficient for evaluating the outputs of next-generation video generation models, which are capable of producing longer and more complex content. Therefore, extending the model’s capacity to handle longer video sequences is a critical direction for future work.
Leverage of Foundation Model. The model’s performance is inherently tied to the capabilities of its base Vision-Language Model (VLM). Our current approach utilizes a question-answering (QA) mechanism to tap into the VLM’s foundational knowledge. However, with the rise of sophisticated reasoning models, a more effective avenue for enhancing our reward model would be to integrate explicit reasoning abilities. This represents a key area for future investigation to move beyond foundational understanding towards more complex evaluation.
@@ 13 More Details of MonetBench
Type Image Video
Content People, Objects, Animals, Story, Human Activity,
Architecture, Landscape, Artificial Scene, Others,
Vehicles, Plants, Food, Natural Animal Activity,
Others, Scenes Physical Phenomena
Challenge Unreal, Style, History, Material, Angle and Lens,
Fine-grained Detail, Emotional Expression,
Color, Famous Character, Color/Tone, Surreal,
Normal, Famous Places, World Knowledge,
Writing, Complex Combo, Special Effects, Text,
Positional, Counting, Spatial Relationship,
Camera Movement,
Logical Consistency,
Style, Temporal Speed
Table 24: Content and Challenge of MonetBench.
Image-MonetBench Construction. We first establish our dataset foundation by collecting 4,038 seed prompts through strategic sampling from established datasets (1,000 from ImageRewardDB (xu2023imagereward), 1,000 from HPDv2 (xu2023imagereward), and 2,038 from Pick-a-Pic (kirstain2023pick)). Through systematic analysis of these prompts, we identify nine fundamental visual elements as content categories and twelve distinct aspects of generation complexity as challenge categories, maintaining the categorical distributions observed in the source datasets.
Video-MonetBench Construction. For video prompt evaluation, we initially sample 20,000 prompts from the VproM (wang2024vidprom) dataset, which are filtered to 13,342 valid entries after removing duplicates and invalid content. Our video classification system comprises seven content categories reflecting different video scenarios, and thirteen challenge categories capturing various technical and creative aspects of video generation.
Prompt Generation and Filtering. To ensure benchmark quality and diversity, we employ ChatGLM (glm2024chatglmfamilylargelanguage) to generate 1,000 new prompts for each benchmark following the established category distributions. Each generated prompt undergoes a three-stage filtering process: (1) Rouge-L (lin-2004-rouge) similarity checking for textual diversity, (2) semantic filtering with a cosine similarity threshold of 0.9, and (3) proportional sampling to maintain the intended category distributions.
Table 24 shows content and challenge categories for MonetBench. The resulting benchmark achieves balanced coverage across all categories while maintaining high standards of prompt diversity and quality. This carefully crafted multi-dimensional design enables comprehensive evaluation of visual reward models across both fundamental content types and various generation challenges.
@@
Details of Classification Proportions. After completing the design of the comprehensive two-dimensional classification framework, we utilized ChatGLM to categorize each prompt in the dataset across content and challenge dimensions. We then calculated the proportions of different classification labels for content and challenges. The content and challenge categories and their respective examples are summarized in Tables 25 and 26. Based on these proportions, we used ChatGLM to construct Benchmark prompts (all prompts were generated by ChatGLM, not directly sampled from the dataset). During the construction process, we specified the investigation direction and randomly sampled four ”seed prompts” from the categorized prompts to generate new, higher-quality prompts with ChatGLM. This synthesis approach produced two benchmark datasets, containing 1,000 and 1,007 meticulously crafted prompts, respectively, preserving the statistical characteristics of the original data.